OpenStudy (anonymous):

How many like terms are in the expression: 10y + 10z - 13x - 4z + 10 + 2x y? 2 3 4 5

OpenStudy (anonymous):

@phi can u help me?

OpenStudy (phi):

this is easier if you use math instead of English

OpenStudy (phi):

like terms means "same variable to the same power"

OpenStudy (anonymous):

@phi yes I understand that

OpenStudy (skullpatrol):

term: In the sum a + b a and b are called terms

OpenStudy (anonymous):

@skullpatrol mmhm I know

OpenStudy (phi):

can you write out the problem using algebra ?

OpenStudy (skullpatrol):

So count them, if you know!

OpenStudy (anonymous):

@phi what do you mean

OpenStudy (phi):

now how many terms have the same variable ?

OpenStudy (skullpatrol):

10z - 4z = (10-4)z = 6z

OpenStudy (anonymous):

@skullpatrol ok @phi 3 or 2

OpenStudy (phi):

only 10z and -4z are "like terms"

OpenStudy (anonymous):

@phi so 2

OpenStudy (phi):

yes, the answer is 2 notice that 10y and 2x y are NOT like terms, because y and x y are not the same in fact x and x*x (or x^2) are not the same (like term)

OpenStudy (phi):

It is good to get this idea straight, because you use it in algebra to simplify (by combining like terms). You need to know like terms when you see them, otherwise the next steps in algebra will get confusing.

OpenStudy (skullpatrol):

similar terms: Two monomials that are exactly alike or are the same except for their numerical coefficients, also called like terms.

OpenStudy (phi):

the one thing that might be confusing is that x*y (written x y) and y*x (written y x) are "like terms" because the order of multiplying does not matter. Normally we put the variables in alphabetical order to keep things easier to see

OpenStudy (skullpatrol):

Recall that a monomial is a numeral, a variable or the product of a numeral and one or more variables.

OpenStudy (skullpatrol):

$$ab=ba$$